OpenStudy (anonymous):

find f'(a) of radical (1-2x) how do i solve this?

OpenStudy (campbell_st):

so \[y = \sqrt{1 -2x}\] or \[y = (1 -2x)^{1/2}\] use the chain rule let u = 1-2x du/dx = -2 so \[y = u^{1/2}\] dy/dx = dy/du x du/dx \[y' = 1/2(1 -2x)^{-1/2} \times -2\] \[y' = -1/(\sqrt{1-2x})\]